Assume that one numver from 1 to 7 is equally likely to be selected at random. Each number corresponds to one of the 7 figures shown
1-circle
2-triangle
3-square
4-square
5-circle
6-triangle
7-circle
determine the probability of selecting a circle, given a number greater than or equal to 5 is selected

Answer by stanbon(75887)

Assume that one numver from 1 to 7 is equally likely to be selected at random. Each number corresponds to one of the 7 figures shown
1-circle
2-triangle
3-square
4-square
5-circle
6-triangle
7-circle
determine the probability of selecting a circle, given a number greater than or equal to 5 is selected
--------------------
P(circle | n >=5) = P(circle AND n>=5)/P(n>=5) = (2/7)/(3/7) = 2/3
====================================================================
Cheers,
Stan H.